OpenStudy (anonymous):

can you factor x^2-4?

OpenStudy (raden):

difference square formula : a^2 - b^2 = (a+b)(a-b)

OpenStudy (raden):

first, x^2-4 = x^2 - 2^2 so, by using formula above it can be (x+2)(x-2)